OpenStudy (anonymous):

HELP ?? Using the properties of exponents and radicals, design at least three different equivalent forms of x². You must show how each one can be simplified back to x² in two or more steps. Stretch your mind and get creative! Keep in mind that something too simple, like x • x would not be acceptable since it takes only one step to convert it to x².

OpenStudy (dumbcow):

use this property \[\large \sqrt{x^{n}} = x^{n/2}\] and \[\large \sqrt[m]{x^{n}} = x^{n/m}\] so \[\large \sqrt{x^{2}}*\sqrt{x^{2}} = x^{2}\] \[\large \sqrt[4]{x^{4}}*\sqrt[4]{x^{4}} = x^{2}\]

OpenStudy (anonymous):

\[(\sqrt[3]{x^2})^(4/3)\] \[\sqrt{\sqrt{\sqrt{x^16}}}\] \[(-3x-4+4x+4)^2\]

OpenStudy (anonymous):

You both put something different ... D: